OpenStudy (anonymous):

simplify x^4-3x^3 over x^2-9

OpenStudy (anonymous):

\[\frac{ x^4-3x^3 }{ x^2-3^2 }=\frac{ x^3(x-3) }{ (x+3)(x-3) }=\frac{ x^3 }{ x+3 }\]

OpenStudy (anonymous):

hope that helps you

OpenStudy (anonymous):

yes thank you i had factored it wrong

OpenStudy (anonymous):

its alrite ;)